# Integrated Information Decomposition Unveils Major Structural Traits of In Silico and In Vitro Neuronal Networks

### v1

@@ Appendix A Experimental methods
[…]
@@ Appendix B Numerical simulation
In short, $N\approx 2800$ neurons ($80\%$ excitatory) were placed uniformly on a circular area with radius $r=1.5~{}\text{mm}$. Following, from each neuron axon growth was simulated by concatenating line segments with randomly fluctuating angles. Once the axon of a neuron $i$ extended within a radius close to an other neuron $j$, a connection from $i$ to $j$ was formed with a $50\%$ probability. The bands obstacles influenced the growth of axons by allowing line segments to cross the obstacle borders with a low probability, and with the complementary probability to redirect the line segment to remain in the originating band.
Neuron dynamics were modelled as Izhikevich regular spiking neurons Izhikevich (2003), supplemented with synaptic dynamics:
@@ Appendix C Calculation of graph measures
The graph theoretic measures used in Figs. 4 and 5 were defined as follows.
@@ C.1 In-degree
The in-degrees $k_{in}^{i}$ was calculated by summing the number of existing directed links toward a node. For a node $i$,
$k_{in}^{i}=|\{j\in V~{}|~{}(j,i)\in A~{}\}|,$
where $(j,i)$ means a directed edge from $j$ to $i$, $V$ is the set of nodes in the network, and $A$ is the set of directed edges in the network.
@@ C.2 Global efficiency $G_{E}$
The global efficiency $G_{E}$ is a measure of the efficiency of information routing through the whole network on average, and was defined as the average of the inverse of the pair-wise network distances Rubinov and Sporns (2010):
$G_{E}=\frac{1}{N(N-1)}\sum_{(i,j)\in A}\frac{1}{d(i,j)},$
where $d(i,j)$ is the shortest geodesic path connecting $i$ to $j$ with $(i,j)\notin A\implies d(i,j)=\infty$, and $N$ denotes the number of neurons.
@@ C.3 Modularity Q
The modularity Q is a measure of the tendency of neurons to connect with other neurons within a group or community rather than with neurons in the rest of the network Rubinov and Sporns (2010). It wass calculated as
$Q=\frac{1}{2m}\sum_{i,j\in V}\left(a_{ij}-\frac{k_{i}k_{j}}{2m}\right)\delta(c_{i},c_{j}),$
where $V$ is the set of neurons, $a_{ij}=1$ if a connection from $j$ to $i$ exists and zero otherwise, $N$ is the number of neurons, $m=(1/2)|a|$, $k_{i}$ the degree of neuron $i$:
$k_{i}=|\{~{}j\in V~{}|~{}(j,i)\in A~{}\lor~{}(i,j)\in A\}|,$
$c_{i}$ denotes the community that neuron $i$ belongs to, and $\delta(\cdot,\cdot)$ is the Kronecker delta such that
@@
Communities were detected using the Louvain algorithm Blondel et al. (2008).

### v2

@@ Appendix A Experimental methods
Neuronal primary cultures were prepared from rat embryonic cortical tissue following the protocol described in Montalà-Flaquer et al. Montalà-Flaquer et al. (2022). Briefly, cortical tissue was dissociated by repeated pipetting and neurons seeded on a polydimethylsiloxane (PDMS) circular surface $6$ mm in diameter that contained topographical motif shaped as parallel bands $300$ $\mu$m wide, $70$ $\mu$m high, and separated by $300$ $\mu$m. Neurons were seeded at day in vitro (DIV) 0 and transduced with the genetically encoded calcium indicator GCaMP6s at DIV 1, and were incubated at 37C, 95% humidity and 5% CO2 for about 2 weeks with periodic culture medium replacements. Cultures contained 80% excitatory neurons and 20% inhibitory ones in their nature stage. Spontaneous activity was recorded through calcium fluorescence imaging in an inverted microscope in combination with a high speed camera that captured images at 20 ms intervals with a spatial resolution of $5.9$ $\mu$m/pixel.
Acquired images were analyzed with the software NETCAL Orlandi et al. (2017) to select regions of interest (ROIs), extract their fluorescent signal as a function of time and identify peaks in fluorescence that revealed neuronal spiking events. A total of 900 ROIs were considered in the experiments, which covered in a grid-like manner the area of the culture Montalà-Flaquer et al. (2022). ROIs were used instead of single neurons since the large size of the cultures in combination with limitations in camera resolution, made not possible to resolve single cell bodies. An ROI contained about $3-5$ neurons, and 900 ROIs were used to balance spatial details and a reasonable analysis time. The trains of identified spikes, extended to all regions of interest, shaped the raw data from which Transfer Entropy was computed Stetter et al. (2012); Montalà-Flaquer et al. (2022).
@@ Appendix B Numerical simulation
The connectivity used in the numerical simulations was established as follows. The $N\approx 2800$ neurons (of which $80\%$ were excitatory) were distributed uniformly on a circular area with radius $r=1.5~{}\text{mm}$. Following, axon growth was simulated to determine the projections made by each neuron. From the center of a neuron $i$, line-segments were concatenated until the axon reached a certain length $L_{i}$. The axon lengths $L_{i}$ were drawn from a Rayleigh distribution with scale parameter $\sigma_{L}=\sqrt{2/\pi}$, such that the average axon length was $\langle L\rangle=1~{}\text{mm}$. Each consecutive line-segment was placed starting from the end of the previous segment with a random deviation in angle of $\phi=0.1~{}\text{radians}$. Once the axon of a neuron $i$ grew into an area of radius $150~{}\mu\text{m}$ around another neuron $j$, a connection from $i$ to $j$ was formed with a $50\%$ probability. The band obstacles influenced the growth of axons by allowing line segments to cross the obstacle borders with a low probability, and with the complementary probability to redirect the line segment to remain in the originating band.
Neuron dynamics were modelled as Izhikevich regular spiking neurons Izhikevich (2003),
$\frac{dV_{i}}{dt}$ $=0.04V_{i}^{2}+5V_{i}+140-U_{i}+\sum_{j=0}^{N}w_{ij}P_{j}+\sigma\eta_{i}$ (12)
$\frac{dU_{i}}{dt}$ $=0.02(0.2V_{i}-U_{i}),$
where $V_{i}$ is the membrane potential, $U_{i}$ a recovery variable, and $\eta_{i}$ a Gaussian white-noise, $\langle\eta_{i}(t),\eta_{i}(\tau)\rangle=\delta(t-\tau)$. The weighted and directed connectivity matrix $W=\{w_{ij}\}$ resulted from the network-growth algorithm described above. The neurons were coupled through the variables $P_{i}$, which evolve following
@@ Appendix C Calculation of graph measures
The graph theoretic measures used in Figs. 4 and 5 were computed using the Brain Connectivity Toolbox Rubinov and Sporns (2010). For in vitro data, these measures were computed on the effective connectivity matrices $A=\{a_{ij}\}$, while for in silico data, these measures were computed on either $W=\{w_{ij}\}$ (ground truth topology) or $A=\{a_{ij}\}$. For clarity, the matrix $A$ was used in the definitions below.
@@ C.1 In-degree
The distribution of in-degrees $p(k_{in})$ in effective networks captured the probability to observe a node in the network (neurons or ROIs) with $k_{in}$ incoming links. For a given node $i$, its number of incoming links was calculated by summing up the number of existing directed links toward the node, as
$k_{in}^{i}=\sum_{j}a_{ij},$ (14)
where $a_{ij}$ are the elements of the effective connectivity matrix $A$, i.e. the set of directed links in the network, with $j\rightarrow i$ indicating that there exist an effective link from neuron $j$ to $i$.
@@ C.2 Global efficiency $G_{E}$
The global efficiency $G_{E}$ quantified the easiness for network-wide integration, i.e., the easiness for a node in the network to topologically reach any other node. $G_{E}$ was defined as the average of the inverse of the pair-wise topological network distances Rubinov and Sporns (2010), as
$G_{E}=\frac{1}{N(N-1)}\sum_{(i,j)\in A}\frac{1}{d(i,j)},$ (15)
where $d(i,j)$ is the shortest geodesic path connecting $i$ to $j$ in the matrix $A$, with $(i,j)\notin A\implies d(i,j)=\infty$, and $N$ denotes the number of neurons. A value $G_{E}=0$ reflected a network in which all nodes were disconnected, while $G_{E}=1$ reflected a complete graph, i.e., a network in which any given node connected to all its possible $N-1$ neighbors.
$G_{E}$ was used as a network metric to characterize the topological organization of the reconstructed effective networks in a quantitative manner and for different thresholds, since its value intuitively accounts for the density of links in a network.
@@ C.3 Modularity Q
The modularity Q measured the tendency of neurons to connect with other neurons within a group or community rather than with other neurons in the rest of the network Rubinov and Sporns (2010). It was calculated as
$Q=\frac{1}{2m}\sum_{i,j\in M}\left(a_{ij}-\frac{k_{i}k_{j}}{2m}\right)\delta(c_{i},c_{j}),$ (16)
where $M=\{1,...,N\}$ is the set of neurons in the network, $a_{ij}$ are the elements of the effective connectivity matrix $A$, $m=(1/2)|a|$, $k_{i}=\sum_{j}a_{ij}+\sum_{j}a_{ji}$ is the total degree ($k^{i}=k^{i}_{in}+k^{i}_{out}$) of neuron $i$, $c_{i}$ denotes the community that neuron $i$ belongs to, and $\delta(\cdot,\cdot)$ is the Kronecker delta such that
@@
Communities were detected using the Louvain algorithm Blondel et al. (2008). $Q$ was used throughout the analysis as a network measure to quantify the tendency of nodes (neurons or ROIs) to organize into communities, with $Q=0$ indicating that the entire network was the only community and $Q=1$ indicating that all nodes were isolated and formed by themselves a community. The interest of using $Q$ and the Louvain algorithm was also based on the fact that they reflected well the organizational features of the in silico networks, in which neurons connected along preferred directions with clear communities, and therefore were considered adequate to later compare the characteristics of reconstructed in silico and in vitro data. However, the used modularity measure could not work adequately with other data, and other definitions could therefore be considered Diez et al. (2015); MacMahon and Garlaschelli (2015); Newman (2016)